# Many-body Fu-Kane-Mele index

### v1

@@ 2.7 The many-body ${\mathbb{Z}}_{2}$-index
With these definitions, we can now state our main theorem which posits the existence of a ${\mathbb{Z}}_{2}$-valued index associated with any symmetric stably SRE state. The empty and spin Hall states described in Section 2.5 give examples of symmetric stably SRE states for which the index takes its two possible values. The index is multiplicative under stacking, and it is an invariant under the action of symmetry preserving LGAs and therefore constant on SPT phases. The theorem thus shows that there are two distinct such phases.
@@
    If $\omega_{SH}$ is a spin Hall state, then $\mathrm{Ind}_{2}(\omega_{\mathrm{SH}})=(-1)^{\mathrm{Ch}(\hat{\mathscr{P}}^{\uparrow})}$.
@@
While we prove that our index generalizes the Fu-Kane-Mele index of the translation-invariant quantum spin Hall states, the same methods of Section 5 could be used to extend the equivalence to translation-invariant cases without spin conservation since the K-theory classification is known to be complete [38, 16].
@@ 3.1.2 Symmetric parent Hamiltonian for $\omega$
Since $\omega_{0}$ is a homogeneous product state there is a 0-chain $H^{(0)}$ such that $H^{(0)}_{x}\in{\mathcal{A}}_{\{x\}}$ is even for each $x\in{\mathbb{Z}}^{2}$ and $\omega_{0}$ is the unique gapped ground state of $H^{(0)}$ with gap one. Since $\omega=\omega_{0}\circ\alpha$ we claim that $\omega$ is the unique gapped ground state of the 0-chain $H^{\prime}$ with gap one, where $H^{\prime}_{x}:=\alpha^{-1}(H^{(0)}_{x})$. Indeed, since $\alpha(A)\in{\mathcal{A}}^{\mathrm{al}}$, the series $\sum_{x\in{\mathbb{Z}}^{2}}[H^{(0)}_{x},\alpha(A)]$ converges in norm, and therefore $[H^{\prime},A]=\sum_{x\in{\mathbb{Z}}^{2}}[\alpha^{-1}(H^{(0)}_{x}),A]=\sum_{x\in{\mathbb{Z}}^{2}}\alpha([H^{(0)}_{x},\alpha(A)])$ is also convergent since $\alpha$ is an automorphism. But then $\omega(A^{*}[H^{\prime},A])=\omega_{0}(\alpha(A)^{*}[H^{(0)},\alpha(A)])\geq 0$ for all $A\in{\mathcal{A}}^{\mathrm{loc}}$ by the remark after Definition 3.1. Hence $\omega$ is a ground state of $H^{\prime}$. If $H^{\prime}$ has another ground state $\omega^{\prime}$, then $\omega^{\prime}\circ\alpha^{-1}$ would be a ground state of $H_{0}$, and so $\omega$ is the unique ground state of $H^{\prime}$. That it is gapped follows by a similar argument. Note that the $H^{\prime}_{x}$ are still even, but they may not be symmetric since $\alpha$ is in general not symmetric. Therefore, we average $H^{\prime}_{x}$ over the time reversal and $U(1)$ symmetries to obtain a symmetric 0-chain $H$:
@@ 3.2.1 Almost local equivalence
see e.g. [8, Lemma 5.4]. If we denote by $-l$ the half-line $-l=(a,-v)$, we have that $K^{l,H}-K^{h,H}=K^{-l,H}$ which satisfies the assumptions of Lemma 3.5, hence so does the TDI (3.31). The Lemma (with $\Lambda$ replacing $\Lambda^{c}$ since it is applied to the cone $\Lambda^{c}$) shows that
@@ 3.2.2 GNS representation and time reversal
Let $(\Pi,{\mathcal{H}},\Omega^{+})$ be the GNS triple of $\omega^{+}$ and fix a homogeneous unitary $U\in{\mathcal{A}}^{\mathrm{al}}$ such that $\omega^{+}=\omega^{-}\circ^{\prime}Ad[U]$, which exists be Proposition 3.8. Then the unit vector $\Omega^{-}:=\Pi(U)\,\Omega^{+}$ represents the state $\omega^{-}$:
@@ 3.3 Definition of the index
Before we study the properties of the index we have just defined, we first show that it has an expression that does not rely on the GNS representation at all. While this expression loses it transparency to physical interpretation, this more ‘algebraic’ formulation hints at the announced invariance under symmetry preserving LGAs. Both formulations can be used in the proofs of the next section but we have chosen to consistency to rely only on Definition 3.15.
@@ 4.1.3 Independence of the half-line
Let $\Gamma\subset{\mathbb{Z}}^{2}$. If $\psi$ is a Kramers singlet or belongs to a Kramers pair then the state $\psi^{\prime}=\psi\circ\rho_{\pi}^{\Gamma}$ is a Kramers singlet if $\psi$ is, and $\psi^{\prime}$ belongs to a Kramers pair of $\psi$ did.
@@ 4.2 Stacking
Let $(\omega_{1},{\mathcal{A}}_{1},\tau_{1})$ and $(\omega_{2},{\mathcal{A}}_{2},\tau_{2})$ be symmetric SRE states. Then $\omega_{1}\,\hat{\otimes}\,\omega_{2}$ is a symmetric SRE state on the fermion system with time-reversal $({\mathcal{A}}_{1}\,\hat{\otimes}\,{\mathcal{A}}_{2},\tau_{1}\,\hat{\otimes}\,\tau_{2})$ and
@@
Let $(\Pi_{i},{\mathcal{H}}_{i},\Omega_{i}^{+})$ be the GNS representation of $\omega_{i}^{+}$ and let $T_{i}$ be the antiunitary that implements time-reversal in the representation $(\Pi_{i},{\mathcal{H}}_{i})$. The index $\mathrm{Ind}_{2}(\omega_{i})$ is defined by $T_{i}^{2}\Omega_{i}^{+}=\mathrm{Ind}_{2}(\omega_{i})\,\Omega_{i}^{+}$, see Lemma 3.14.
@@ 4.3 Invariance under symmetry preserving LGAs
Let $\omega$ a symmetric SRE state and let $G$ be a symmetry preserving TDI. Then
@@ 4.4 Extension of the index to symmetric stably SRE states
5 Free fermion examples
In this section, we establish that the quasi-free spin Hall states $\omega_{\mathrm{SH}}$ described in Example 2.5.2 are stably SRE and have a non-trivial value of the index, as claimed in Theorem 2.5.
5.1 Proof that spin Hall states are stably SRE
We will first prove a more general fact and then specialise to the setting of Lemma 2.3.
Lemma 5.1.
Proof.
Since isomorphism classes of complex vector bundles over the torus are completely classified by the Chern number, our assumption $\mathrm{Ch}(\hat{\mathscr{P}})=0$ implies that the bundle $E_{\hat{\mathscr{P}}}$ is isomorphic to a trivial bundle. It now follows from [13, Theorem 23.10], and noting that the torus has a good open cover of 7 sets ([36, Theorem 5.3]), that there is a homotopy $\hat{\mathscr{P}}:[0,1]\times{\mathbb{T}}\rightarrow\mathrm{Gr}_{r}({\mathbb{C}}^{n})$ such that $\hat{\mathscr{P}}_{0}=\hat{P}$ and $\hat{\mathscr{P}}_{1}$ is constant. By [41, Theorem 10.22] we can moreover take this homotopy to be smooth. ∎
$({\mathscr{A}}f)(x)=\frac{1}{4\pi^{2}}\int_{{\mathbb{T}}}\,{\mathrm{d}}k\,\mathrm{e}^{{\mathrm{i}}k\cdot x}\hat{\mathscr{A}}(k)\hat{f}(k)$ (5.1)
$|\langle\delta_{x}\otimes u,{\mathscr{A}}\,\delta_{y}\otimes v\rangle|\leq g_{A}(\mathrm{dist}(x,y))$ (5.2)
Lemma 5.2.
Let
$\hat{\mathscr{P}}:[0,1]\times{\mathbb{T}}\rightarrow\mathrm{Proj}_{r}({\mathbb{C}}^{n}):(t,k)\mapsto\hat{\mathscr{P}}_{t}(k)$ (5.3)
${\mathrm{i}}\frac{{\mathrm{d}}{\mathscr{V}}_{t}}{{\mathrm{d}}t}={\mathscr{G}}_{t}{\mathscr{V}}_{t},\qquad{\mathscr{V}}_{0}=\operatorname{\mathbbm{1}},\qquad\text{and}\quad{\mathscr{P}}_{t}={\mathscr{V}}_{t}{\mathscr{P}}_{0}{\mathscr{V}}_{t}^{*}.$ (5.4)
The generators ${\mathscr{G}}_{t}$ have super-polynomial off-diagonal decay in the sense of (5.2), where $g\in{\mathcal{F}}$ is independent of $t$.
Proof.
Let
$\hat{\mathscr{G}}_{t}(k)={\mathrm{i}}\left[\frac{\partial}{\partial t}\hat{\mathscr{P}}_{t}(k),\hat{\mathscr{P}}_{t}(k)\right]$ (5.5)
${\mathrm{i}}\frac{\partial}{\partial t}\hat{\mathscr{V}}_{t}(k)=\hat{\mathscr{G}}_{t}(k)\hat{\mathscr{V}}_{t}(k),\qquad\hat{\mathscr{V}}_{0}(k)=\operatorname{\mathbbm{1}}.$ (5.6)
Then
$\hat{\mathscr{P}}_{t}(k)=\hat{\mathscr{V}}_{t}(k)\hat{\mathscr{P}}_{0}(k)\hat{\mathscr{V}}_{t}(k)^{*}.$ (5.7)
Lemma 5.3.
Let $[0,1]\ni t\mapsto{\mathscr{G}}_{t}$ be a norm-continuous family of self-adjoint operators on $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{n})$ such that (5.2) holds uniformly in $t$. Let ${\mathscr{V}}_{t},{\mathscr{P}}_{t}$ be the corresponding one-parameter family of unitaries given by (5.4). Let ${\mathcal{A}}$ be the CAR algebra over $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{n})$ and let $\omega_{t}=\omega_{{\mathscr{P}}_{t}}$ be the gauge-invariant quasi-free state on ${\mathcal{A}}$ corresponding to the projection ${\mathscr{P}}_{t}$, see (2.11). Then there is a $U(1)$ symmetry preserving TDI $F$ such that $\omega_{t}=\omega_{0}\circ\alpha_{t}^{F}$ for all $t\in[0,1]$.
Proof.
${\mathscr{G}}_{t}(x,i;y,j):=\langle\delta_{x}\otimes e_{i},\,{\mathscr{G}}_{t}\,\delta_{y}\otimes e_{j}\rangle$ (5.8)
$F_{x}(t)=\frac{1}{2}\sum_{y\in{\mathbb{Z}}^{2}}\sum_{i,j=1}^{n}\,\left({\mathscr{G}}_{t}(x,i;y,j)a^{*}_{x,i}a_{y,j}+\overline{{\mathscr{G}}_{t}(x,i;y,j)}a^{*}_{y,j}a_{x,i}\right)$ (5.9)
$\sum_{x\in{\mathbb{Z}}^{2}}{\mathrm{i}}\left[F_{x}(t),a(f)\right]=a({\mathrm{i}}{\mathscr{G}}_{t}f)$ (5.10)
which implies that
$\alpha^{F}_{t}(a(f))=a({\mathscr{V}}_{t}^{*}f).$ (5.11)
for any $f\in l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{n})$.
With this, we check that
$(\omega_{0}\circ\alpha^{F}_{t})\big{(}a^{*}(f)a(g)\big{)}=\omega_{0}\big{(}a^{*}({\mathscr{V}}_{t}^{*}f)a({\mathscr{V}}_{t}^{*}g)\big{)}=\left\langle{\mathscr{V}}_{t}^{*}g,{\mathscr{P}}_{0}\,{\mathscr{V}}_{t}^{*}f\right\rangle=\left\langle g,{\mathscr{P}}_{t}\,f\right\rangle=\omega_{t}\big{(}a^{*}(g)a(f)\big{)}.$ (5.12)
We are now equipped to prove Lemma 2.3. We briefly recall the setting described in Section 2.5.2. The one particle Hilbert space is ${\mathcal{K}}_{2}=L^{2}({\mathbb{T}};{\mathbb{C}}^{2})\oplus L^{2}({\mathbb{T}};{\mathbb{C}}^{2})$ and the state $\omega_{\mathrm{SH}}$ corresponds to the projection ${\mathscr{P}}_{\mathrm{SH}}={\mathscr{P}}^{\uparrow}\oplus{\mathscr{P}}^{\downarrow}$ whose Fourier transform is a smooth projection valued map of the form
$\hat{\mathscr{P}}_{\mathrm{SH}}(k)=\hat{\mathscr{P}}^{\uparrow}(k)\oplus\hat{\mathscr{P}}^{\downarrow}(k)$ (5.13)
Here $\hat{\mathscr{P}}^{\downarrow}(k)$ is the complex conjugate of $\hat{\mathscr{P}}^{\uparrow}(k)$. It follows that the projection valued maps $\hat{\mathscr{P}}^{\uparrow}$ and $\hat{\mathscr{P}}^{\downarrow}$ have opposite Chern numbers, and since the Chern number is additive under direct sums of bundles, this implies $\mathrm{Ch}(\hat{\mathscr{P}}_{\mathrm{SH}})=\mathrm{Ch}(\hat{\mathscr{P}}^{\uparrow})+\mathrm{Ch}(\hat{\mathscr{P}}^{\downarrow})=0$.
Proof of Lemma 2.3.
In order to apply Lemma 5.1, we expand the ‘ambient dimension’ from $4$ to $14$ by considering the algebra ${\mathcal{A}}^{\prime}$ over $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{4})\oplus l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{10})$ and the symmetric state $\omega^{\prime}=\omega_{\mathrm{SH}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ corresponding to the projection ${\mathscr{P}}={\mathscr{P}}_{\mathrm{SH}}\oplus 0$ and consequently $\hat{\mathscr{P}}=\hat{\mathscr{P}}_{\mathrm{SH}}\oplus 0$. The projection-valued map $\hat{\mathscr{P}}=\hat{\mathscr{P}}_{\mathrm{SH}}\oplus 0$ takes values in $\mathrm{Proj}_{2}({\mathbb{C}}^{14})$ and has $\mathrm{Ch}(\hat{\mathscr{P}})=0$. Now Lemma 5.1 provides a smooth homotopy $(t,k)\mapsto\hat{\mathscr{P}}_{t}(k)\in\mathrm{Proj}_{2}({\mathbb{C}}^{12})$ interpolating between $\hat{\mathscr{P}}$ and a constant projection.
Let ${\mathscr{P}}_{t}$ be the inverse Fourier transform of $\hat{\mathscr{P}}_{t}$. We then apply Lemma 5.2 to obtain unitaries ${\mathscr{V}}_{t}$ and generators ${\mathscr{G}}_{t}$ satisfying the assumption of Lemma 5.3. The latter lemma now provides a $U(1)$-invariant TDI $F$ such that $\omega_{t}^{\prime}=\omega_{0}\circ\alpha_{t}^{F}$ where $\omega_{t}^{\prime}=\omega^{\prime}_{{\mathscr{P}}_{t}}$ is the quasi-free state on ${\mathcal{A}}^{\prime}$ determined by the projection ${\mathscr{P}}_{t}$. This family of states interpolates between $\omega_{\mathrm{SH}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ and $\omega_{{\mathscr{P}}_{1}\oplus 0}$. Since $\hat{\mathscr{P}}_{1}$ is constant, its Fourier transform ${\mathscr{P}}_{1}$ is strictly local and $\omega_{1}=\omega_{{\mathscr{P}}_{1}\oplus 0}$ is a product state. We conclude that $\omega_{\mathrm{SH}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ is LGA equivalent to a product state. Since $\omega_{\mathrm{empty}}$ is also a product state, we conclude that $\omega_{\mathrm{SH}}$ is stably SRE, see Section 2.4. ∎
5.2 Index of spin Hall states
Let $\omega_{\mathrm{SH}}$ be a spin Hall state on the CAR algebra ${\mathcal{A}}$ over $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{4})$. We have just proved that $\omega_{\mathrm{SH}}$ is stably SRE; Concretely, $\omega=\omega_{\mathrm{SH}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ is a symmetric SRE state and it corresponds to the Fermi projection ${\mathscr{P}}={\mathscr{P}}_{\mathrm{SH}}\oplus 0$ on $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{4})\oplus l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{10})$ where ${\mathscr{P}}_{\mathrm{SH}}={\mathscr{P}}^{\uparrow}\oplus{\mathscr{P}}^{\downarrow}$ is the Fermi projection of the spin Hall state.

### v2

@@ 2.7 The many-body ${\mathbb{Z}}_{2}$-index
With these definitions, we can now state our main theorem which posits the existence of a ${\mathbb{Z}}_{2}$-valued index associated with any symmetric stably SRE state. The $\operatorname{\mathrm{AII}}$ states described in Section 2.5 give examples of symmetric stably SRE states for which the index takes its two possible values. The index is multiplicative under stacking, and it is an invariant under the action of symmetry preserving LGAs and therefore constant on SPT phases. The theorem thus shows that there are two distinct such phases.
@@
    If $\omega_{{\mathscr{P}}}$ is an $\operatorname{\mathrm{AII}}$ state that is stably SRE, then $\mathrm{Ind}_{2}(\omega_{{\mathscr{P}}})=\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})$.
@@
The first point of this Theorem says that our index generalizes the Fu-Kane-Mele index of $\operatorname{\mathrm{AII}}$ states that are stably SRE. While we expect that all $\operatorname{\mathrm{AII}}$ states are stably SRE, currently we can only prove it in the translation invariant case:
Proposition 2.5.
Translation invariant $\operatorname{\mathrm{AII}}$ states are stably SRE.
This proposition, which is proven in Appendix B, provides plenty of states for which our index takes its possible values.
@@ 3.1.2 Symmetric parent Hamiltonian for $\omega$
Note that a symmetric parent Hamiltonian is not time reversal preserving as defined in Section 2.3, it is instead time reversal invariant, which is not the same thing because time reversal is an anti-linear symmetry.
@@
Since $\omega_{0}$ is a homogeneous product state there is a 0-chain $H^{(0)}$ such that $H^{(0)}_{x}\in{\mathcal{A}}_{\{x\}}$ is even for each $x\in{\mathbb{Z}}^{2}$ and $\omega_{0}$ is the unique gapped ground state of $H^{(0)}$ with gap one. Since $\omega=\omega_{0}\circ\alpha$ we claim that $\omega$ is the unique gapped ground state of the 0-chain $H^{\prime}$ with gap one, where $H^{\prime}_{x}:=\alpha^{-1}(H^{(0)}_{x})$. Indeed, since $\alpha(A)\in{\mathcal{A}}^{\mathrm{al}}$, the series $\sum_{x\in{\mathbb{Z}}^{2}}[H^{(0)}_{x},\alpha(A)]$ converges in norm, and therefore $[H^{\prime},A]=\sum_{x\in{\mathbb{Z}}^{2}}[\alpha^{-1}(H^{(0)}_{x}),A]=\sum_{x\in{\mathbb{Z}}^{2}}\alpha([H^{(0)}_{x},\alpha(A)])$ is also convergent since $\alpha$ is an automorphism. But then $\omega(A^{*}[H^{\prime},A])=\omega_{0}(\alpha(A)^{*}[H^{(0)},\alpha(A)])\geq 0$ for all $A\in{\mathcal{A}}^{\mathrm{loc}}$ by the remark after Definition 3.1. Hence $\omega$ is a ground state of $H^{\prime}$. If $H^{\prime}$ has another ground state $\omega^{\prime}$, then $\omega^{\prime}\circ\alpha^{-1}$ would be a ground state of $H_{0}$, and so $\omega$ is the unique ground state of $H^{\prime}$. That it is gapped follows by a similar argument. Note that the $H^{\prime}_{x}$ are still even, but they may not be symmetric since $\alpha$ is in general not symmetric. Therefore, we average $H^{\prime}_{x}$ over the time reversal and $U(1)$ symmetries to obtain a symmetric Hamiltonian $H$:
@@ 3.2.1 Almost local equivalence
see e.g. [11, Lemma 5.4]. If we denote by $-l$ the half-line $-l=(a,-v)$, we have that $K^{l,H}-K^{h,H}=K^{-l,H}$ which satisfies the assumptions of Lemma 3.5, hence so does the TDI (3.31). The Lemma (with $\Lambda$ replacing $\Lambda^{c}$ since it is applied to the cone $\Lambda^{c}$) shows that
@@ 3.2.2 GNS representation and time reversal
Let $(\Pi,{\mathcal{H}},\Omega^{+})$ be the GNS triple of $\omega^{+}$ and fix a homogeneous unitary $U\in{\mathcal{A}}^{\mathrm{al}}$ such that $\omega^{+}=\omega^{-}\circ\mathrm{Ad}[U]$, which exists be Proposition 3.8. Then the unit vector $\Omega^{-}:=\Pi(U)\,\Omega^{+}$ represents the state $\omega^{-}$:
@@ 3.3 Definition of the index
Before we study the properties of the index we have just defined, we first show that it has an expression that does not rely on the GNS representation at all. While this expression loses it transparency to physical interpretation, this more ‘algebraic’ formulation hints at the announced invariance under symmetry preserving LGAs. Both formulations can be used in the proofs of the next section but we have chosen to consistently rely only on Definition 3.15.
@@ 4.1.3 Independence of the half-line
Let $\Gamma\subset{\mathbb{Z}}^{2}$. If $\psi$ is a Kramers singlet or belongs to a Kramers pair then the state $\psi^{\prime}=\psi\circ\rho_{\pi}^{\Gamma}$ is a Kramers singlet if $\psi$ is, and $\psi^{\prime}$ belongs to a Kramers pair if $\psi$ did.
@@ 4.2 Stacking
Let $(\omega_{1},{\mathcal{A}}_{1},\tau_{1})$ and $(\omega_{2},{\mathcal{A}}_{2},\tau_{2})$ be symmetric SRE states. Then $\omega_{1}\,\hat{\otimes}\,\omega_{2}$ is a symmetric SRE state on the fermion system with time reversal $({\mathcal{A}}_{1}\,\hat{\otimes}\,{\mathcal{A}}_{2},\tau_{1}\,\hat{\otimes}\,\tau_{2})$ and
@@
Let $(\Pi_{i},{\mathcal{H}}_{i},\Omega_{i}^{+})$ be the GNS representation of $\omega_{i}^{+}$ and let $T_{i}$ be the antiunitary that implements time reversal in the representation $(\Pi_{i},{\mathcal{H}}_{i})$. The index $\mathrm{Ind}_{2}(\omega_{i})$ is defined by $T_{i}^{2}\Omega_{i}^{+}=\mathrm{Ind}_{2}(\omega_{i})\,\Omega_{i}^{+}$, see Lemma 3.14.
@@ 4.3 Invariance under symmetry preserving LGAs
Let $\omega$ be a symmetric SRE state and let $G$ be a symmetry preserving TDI. Then
@@ 4.4 Extension of the index to symmetric stably SRE states
5 Index of $\operatorname{\mathrm{AII}}$ states
In this section, we establish that stably SRE $\operatorname{\mathrm{AII}}$ states have an index that agrees with the Fu-Kane-Mele invariant, as claimed in Theorem 2.4. We fix a single particle Hilbert space ${\mathcal{K}}_{m}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2m})$ with time reversal ${\mathscr{T}}$ as described in Section 2.5. We consider an orthogonal projector ${\mathscr{P}}$ on ${\mathcal{K}}_{m}$ that is exponentially local (see (2.16)) and time reversal invariant (${\mathscr{T}}{\mathscr{P}}{\mathscr{T}}^{*}={\mathscr{P}}$). We further assume that the corresponding quasi-free state $\omega_{{\mathscr{P}}}$ is stably SRE, i.e. we can stack with an empty product state $\omega_{\mathrm{empty}}$ on the CAR algebra of some ${\mathscr{K}}_{m^{\prime}}$ to obtain an SRE state which is again quasi-free, corresponding to the exponentially local time reversal invariant projector ${\mathscr{P}}\oplus 0$ on ${\mathscr{K}}_{m}\oplus{\mathscr{K}}_{m^{\prime}}$. By definition, the index of $\omega_{{\mathscr{P}}}$ is equal to the index of this SRE state. Since the Fu-Kane-Mele index of ${\mathscr{P}}$ and ${\mathscr{P}}\oplus 0$ are the same, we can without loss of generality assume that $\omega_{{\mathscr{P}}}$ is SRE to begin with.